OpenStudy (hba):

(x+iy)^2=3+4i,then (x,y) is

OpenStudy (hba):

@hartnn

OpenStudy (hba):

Hints

hartnn (hartnn):

(a+b)^2=a^2+2ab+b^2 i^2=-1 these are hints

OpenStudy (hba):

So x^2-y^2+i2xy=3+4i

OpenStudy (hba):

So Now i equate right ? @hartnn

hartnn (hartnn):

yes

OpenStudy (hba):

So x^2-y^2=3 and xy=2

OpenStudy (hba):

This is gonna be long then :( Any Short-Cut @hartnn

hartnn (hartnn):

i don't think there is any shortcut to this....square roots of complex numbers are always difficult to find.....

OpenStudy (hba):

Well I only get 30 seconds per question so it is gonna be diff for me then :(

hartnn (hartnn):

why 30 sec ??

OpenStudy (hba):

Because It Is Multiple choice questions

hartnn (hartnn):

then u must directly put options and try

hartnn (hartnn):

don't go for solving

hartnn (hartnn):

what are the options ?

OpenStudy (hba):

A)(2,1),B(1,2),C(-2,-1),D)A and B , E) A And C

hartnn (hartnn):

check for which x^2-y^2=3

OpenStudy (hba):

k

OpenStudy (hba):

The Answer Is E :)

hartnn (hartnn):

yup

OpenStudy (hba):

thanks alot. :)